# Don’t Trust Generative Agents to Mimic Communication on Social Networks Unless You Benchmarked their Empirical Realism

### v2

@@ Abstract
Simon Münker Trier University muenker@uni-trier.de          Nils Schwager Trier University s2nsschw@uni-trier.de          Achim Rettinger Trier University rettinger@uni-trier.de
@@ 1 Introduction
Across the world, there are ongoing legislative initiatives to regulate social media platforms. For example, Australia and Spain plan to restrict the use of social networks by minors under the age of 16111Australia approves social media ban on under-16s: https://www.bbc.com/news/articles/c89vjj0lxx9o, and TikTok is already banned in several countries due to the spread of misinformation Basch et al. (2021) and harmful content Weimann and Masri (2023). According to the Digital Service Act (DSA) of the EU, providers of very large online platforms need to "identify, analyze and assess any systemic risks […] from the use made of their services" like "any actual or foreseeable negative effects on civic discourse […]."
However, obtaining realistic quantitative evidence concerning the risks that online platforms pose, and specifically what role platform mechanisms, like ranking and recommendation play, is still being debated in computational social and communication science Wang et al. (2016). The key challenge to obtaining robust insights is to produce counterfactual evidence that tests whether alternative network mechanisms would have had a different (more favorable) outcome. This can best be achieved by simulation, since alternative designs can be tested without the involvement of real users.
The validity of the insights obtained by simulation can only be as good as it represents reality. With the advent of Large Language Models (LLMs) for the imitation of persons in social simulation, the complexity of communication in social simulation has become arguably more realistic Chuang et al. (2024) compared to previously used symbolic agent communication languages. Whether LLM-based simulation can replicate social science studies with human participants to a degree sufficient for robust scientific insights needs an experimental design that improves reproducibility and provides a measurement of empirical realism. Testing under which conditions and to what extent agents can mimic human behavior is crucial to validate simulation-based empirical findings.
This paper contributes to the rigor of conducting simulations of social networks with LLMs by:
  • Formalizing social networks to $i)$ quantify and standardize their simulation and ii)ii) allow to benchmark their empirical realism.
  • Providing different approaches and benchmark data sets for the imitation of different types of user communication on $\mathbb{X}$, like posting and replying.
  • Evaluating the empirical realism and identifying key findings about the potential and limitations of mimicking users using LLMs.
@@ 2 Related Work
There is a long history of distributed artificial intelligence research and multi-agent systems in modeling agents to resemble human decision-making processes Weiss (1999). Those traditional approaches define top-down cognitive models of white-box agents with abstract concepts and deductive reasoning processes that control the agent‘s behavior. Similarly, the rational choice model and game theoretic models are concerned with explicit motives and perceived restrictions underlying human behavior. The social sciences have come up with many alternative explanations of human behavior. However, all of those theories tend to be rather inaccurate when it comes to predicting human behavior. In contrast, this paper is solely concerned with mimicking the communication of users as precisely as possible and does not care to explain or control the behavior. This implies that the agents are generating natural language and not symbolic agent languages such as FIPA ACL O’Brien and Nicol (1998).
The usage of LLMs as human simulacra (representation) began with the application as non-player characters (NPCs) in a Sims-style [^2] The Sims is a series of life simulation video games developed by Maxis and published by Electronic Arts game world to simulate interpersonal communication and day-to-day lives Park et al. (2023). The results showed superficially authentic and believable human behavior. Current research interest revolves around improving those agents in a technical sense, by refining prompt schemes and model-internal feedback loops Wang et al. (2024). However, the application of LLMs as synthetic characters has expanded beyond gaming environments to various fields of social science research Argyle et al. (2023). Researchers are increasingly exploring the potential of these models to simulate human participants in studies Chuang et al. (2024); Papachristou and Yuan (2024), particularly in contexts where the obtaining of real-world data is challenging or ethically complex. Those disciplines have already started to use these models as replacements in social studies arguing that conditioning through prompting causes the systems to accurately emulate response distributions from a variety of human subgroups Argyle et al. (2023); Chang et al. (2025).
Although these applications show promise, they also raise significant methodological and ethical questions. The reliability and validity of using LLMs to represent human behavior and cognition is still subject to debate Larooij and Törnberg (2025). Current research raises concerns about potential biases in training data that lead to misrepresentation of certain groups or points of view Abid et al. (2021); Hutchinson et al. (2020). Furthermore, the use of LLMs as replacements for human participants in social research raises ethical considerations about informed consent and the potential for misuse or misinterpretation of results. Without a deeper understanding of the model’s representations of ideologies, we risk oversimplifying complex human behaviors and social dynamics. Especially since these approaches ignore the LLMs lack of embodiment in the physical world Argyle et al. (2023). This disembodied nature means they lack the grounding in physical reality – expressed by cultural contexts, physical environments, and interpersonal relationships – that shapes human cognition, perception, and decision making Hussein (2012).
Although less complex than modeling communicative behavior, research on LLMs generating synthetic public opinion demonstrates whether base LLMs can accurately represent opinions in a sociocultural context. Although LLMs show an excessive ability to replicate US personas - closely approximating real world opinion distributions - they underperform when being tasked with replicating more complex non-English populations Argyle et al. (2023); Ma et al. (2024). Ma et al. demonstrate sensitivity to variable inclusion/exclusion. Similarly Tjuatja et al. report significant behavioral changes in LLM trained with RLHF due to question perturbations, attributing this behavior to the training scheme, stressing the need to adjust LLMs to case-specific data to improve grounding in reality.
@@ 3 Modeling Online Social Networks
In this section, we provide a concise formal framework for modeling Twins of Online Social Networks (TWONs). Both the users (Sect. 3.1) and the network mechanics (Sect. 3.2) are captured. Fig. 1 provides an illustrating example.
Figure 1: Illustrating example: Agent 2 sends a message to agent 1, which updates its model and generates a reply. Agent 2 receives a curated list of all incoming messages in $t+1$ and updates its model based on that.
@@ 3.1 Modeling Agents
$M$ is a set of messages, where $m_{t}^{(i,j)}$ is a message from agent $i$ to agent $j$ at time $t$. Consequently, $m_{t}^{(i,\cdot)}$ indicates a message that agent $i$ transmits to all other agents and $M_{t}^{(\cdot,i)}$ indicates all messages sent to agent ii. [^4] Note, that messages can be of any modality or type (including images, videos, likes, shares,…) Accordingly, $M_{t}^{i}$ indicates all messages that agent $i$ sends and receives at time $t$.
@@
The communicative behavior $b^{i}()$ of the agent $i$ is a function that maps its current state $a_{t}^{i}$ to the messages $m_{t+1}^{(i,j)}$ he is sending next to agents $j$: $b^{i}_{t}(a_{t}^{i})=M_{t}^{(i,j)}$.
@@ 3.2 Modeling Network Mechanics
The mechanics of the communication channel (here, an online social network) is defined by a function $r^{i}_{t}()$ that adapts the agent $i$’ to the incoming messages $M_{t}^{(\cdot,i)}$ to $\widetilde{M}_{t}^{(\cdot,i)}$, returning a manipulated list of messages. Standard manipulations of the list are filtering or ranking, but could also include the manipulation of content555e.g., by adding community notes or fake news warnings or the addition of messages to the list that have not been directly sent between agents666e.g., advertisements, paid content, trending content,…. Note that this function can be personalized can be adapted over time.
@@
4 Simulating Online Social Networks
Based on the formal framework introduced in Sect. 3 we define the required tasks to construct a twin of a real-world online social network by replicating its behavior at the user (Sect. 4) and system level (Sect. 4). Finally, this allows us to check the empirical realism of such TWONs (Sect. 4).
Behavior $b$ Agent AA Task Type Input $\widetilde{M}_{t}^{(\cdot,i)}$ Output $m_{t+1}^{(i,\cdot)}$ Loss $L_{b}$ Metric(s) $q$
@@ 4 Simulating Online Social Networks
The task of machine learning an agent-based simulation of human social media communication behavior at user level is to estimate the function $b^{i}$ of each agent $i$, which given real-world observations of discourses $(M_{0}^{i},M_{1}^{i},...,M_{t}^{i})$ predicts its next message $m_{t+1}^{i}$. Thus, the objective can be formulated as minimizing a loss function
@@
that compares the predicted $\widehat{M}_{t}^{(\cdot,i)}$ with its observed value. [^8] Here, a ranking metric like MRR or NDCG seems best suited.
@@ 5 Case Study Design: Imitating $\mathbb{X}$ Users
In the following sections, we instantiate the formal framework described in the previous sections based on observations from the social network $\mathbb{X}$. However, we acknowledge that our current implementation represents a simplified subset of the full framework, focusing primarily on the user behavior function $b^{i}$ while making simplifying assumptions about other components. Our focus is on estimating the communicative behavior $b^{i}$ of a set of users $A$ from $\mathbb{X}$. We do so by minimizing $L_{b}(\widehat{m}_{t+1})$, where $m$ can be a message of type post, reply, or like.
@@ 6 Fitting Agents to User Data from $\mathbb{X}$
Our experiments are based on two datasets – English and German – collected from $\mathbb{X}$. The sets are collected around keywords concerning the political discourses in the US and Germany during the first half of 2023. The samples contain two types of content: a) Tweets (posts) from delegates of the national parliament concerning political decisions – DE: $155.000$, EN: $930.000$ and b) replies from regular Twitter users towards these decisions or opinions – DE: $185.000$, EN: $17.800.000$. In addition, we labeled the topics of posts and replies using zero-shot prompt classification with Llama 3.1 70B Zheng et al. (2023). This yields two distinct agent types, the a) political actor for posts and the b) unspecified Twitter user for replies, both across two languages.
@@
Retweets: Tweets starting with *"RT"* mark a shared post. The author does not add new content, but replicates the original text and shares it with their followers. We exclude these samples because we cannot verify if the shared content is in line with the retweeting author’s opinion.
@@
We train two model adapters on top of Llama-3.2-3B-Instruct (Dubey et al., 2024). Both models are trained using the supervised fine-tuning (SFT) paradigm through the transformer reinforcement learning (TRL) library using the default pipeline. The trained adapters are unmodified LoRA (Hu et al., 2021) matrices provide by the parameter-efficient fine-tuning (PEFT) package. We optimize for two objectives (posting and replying) with contents from the original users as the labeled data.
@@ 7.1 Evaluation of Imitated Posting and Replying Behavior
An initial observation is that BLEU scores Papineni et al. (2002) in all in-context settings are low. This is also true for the n-gram metrics. However, the fine-tuning-based approach shows significant improvements specifically in the English reply task, but also for uni-gram matches and the overall length ratio. Beyond such token-overlap-metrics, we tested semantic similarity metrics: we embedded all samples using Twitter-fine-tuned BERT (Zhang et al., 2023) and calculated the distance between generations and original samples. Again, the results generated from the fine-tuned model show a significantly shorter distance to its original counterpart compared to the in-context model. The correlation between the predicted emotions in the original and generated samples using a BERT-based multi-label emotion classifier (Barbieri et al., 2020) is less clear. The fine-tuned approach only marginally outperforms the prompt-only approach on the posting task.
@@ 7.4 Imitation Limitations
Our experimental setup has further limitations regarding both data selection and the system’s capacity to generate discourses containing well-structured standpoints and arguments. Although our models demonstrate such capabilities under certain conditions, their performance characteristics primarily reflect the behavioral patterns of the most active users within our dataset, thus mirroring the distinctive communication dynamics observed in the selected Twitter community. This sampling bias raises questions about the generalizability of our findings in different contexts of social networks and user populations.
@@ 8.2 Future Research
Unfortunately, our research objectives, demonstrating "empirical realism" (aka "ecological validity") of our simulation, do not allow us to use synthetic personas rather than real user data. Creating fictional user profiles with diverse characteristics would eliminate privacy concerns, but without demonstrating the validity of the simulation. We argue that you can only claim the applicability of findings produced via social simulations to the real world once you have proven the "empirical realism" (aka "ecological validity") of your simulation. The key component for showing this is replicability on an individual level, which requires real-world data from individuals. Synthetic personas cannot be used to prove "empirical realism", but real-world data are essential to prove the empirical realism of synthetic personas. And once you prove the empirical realism of synthetic personas with real-world data, you again have a sufficiently similar replica of real-world users. Thus, you cannot have both: A valid and provable realistic social simulation without using real-world user data to model agents and measure their empirical realism.

### v3

@@ Abstract
Simon Münker Trier University muenker@uni-trier.de          Nils Schwager Trier University schwager@uni-trier.de          Achim Rettinger Trier University rettinger@uni-trier.de
@@ 1 Introduction
Social media platforms face mounting regulatory pressure worldwide. Australia and Spain are implementing age restrictions for users under 16111Australia approves social media ban on under-16 https://www.bbc.com/news/articles/c89vjj0lxx9o, while several countries have banned TikTok due to concerns about misinformation Basch et al. (2021) and harmful content Weimann and Masri (2023). The EU’s Digital Service Act mandates that large platforms "identify, analyze and assess any systemic risks […] from the use made of their services," including "actual or foreseeable negative effects on civic discourse."
Meeting these regulatory requirements demands robust quantitative evidence about platform risks—particularly the role of algorithmic mechanisms like ranking and recommendation systems. Yet obtaining such evidence remains challenging in computational social and communication science Wang et al. (2016). The fundamental problem is causal inference: determining whether alternative platform designs would produce different, potentially more favorable outcomes. This question is best addressed through simulation, which enables testing counterfactual scenarios without exposing real users to potentially harmful conditions.
The validity of simulation-based insights depends critically on how faithfully simulations represent reality. Recent advances in Large Language Models (LLMs) have enabled more sophisticated agent communication in social simulations Chuang et al. (2024), surpassing the realism of traditional symbolic agent languages. However, this increased sophistication raises a crucial question: Can LLM-based agents replicate human social behavior with sufficient fidelity to produce scientifically robust insights?
Answering this question requires systematic experimental designs that improve reproducibility and provide quantitative measures of empirical realism, the degree to which simulated behaviors match observed human behaviors. Without such validation, simulation findings lack the credibility needed to inform policy decisions or scientific understanding.
Contributions
Our work advances the rigor of LLM-based social network simulation through three key contributions:
Framework
We formalize social networks to (i) standardize their simulation and (ii) enable systematic benchmarking of empirical realism through quantifiable metrics.
Case Study
We instantiate this framework with multiple approaches for imitating user communication on $\mathbb{X}$, including posting, replying, and reply-likelihood prediction tasks across English and German datasets.
Validation
We benchmark empirical realism across tasks, languages, and modeling approaches, revealing when and how LLMs can faithfully replicate user behavior, and crucially, when they cannot.
@@ 2 Related Work
2.1 From Symbolic Agents to Neural Language Models
Social simulation has evolved from symbolic O’Brien and Nicol (1998), rule-based agents to data-driven neural approaches. Traditional multi-agent systems Weiss (1999) employed explicit cognitive models with deductive reasoning, while game-theoretic frameworks relied on rational choice assumptions. These approaches prioritized theoretical interpretability and control over predictive accuracy, a reasonable trade-off when modeling capabilities were limited.
Our work diverges fundamentally from this tradition. Rather than constructing agents from theoretical principles, we aim to replicate observed communication patterns as precisely as possible. This shifts the goal from explanation to imitation, and from symbolic representations to natural language generation. The question is not whether agents follow theoretically motivated rules, but whether they produce outputs statistically indistinguishable from real users.
2.2 LLMs as Human Simulacra: Promise and Limitations
The use of LLMs as human surrogates began with believable non-player characters in game environments Park et al. (2023), demonstrating superficially authentic behavior. This success sparked broader interest in using LLMs as synthetic participants across social sciences Argyle et al. (2023); Chuang et al. (2024); Papachristou and Yuan (2024). Proponents argue that appropriate prompting enables models to accurately emulate response distributions from diverse demographic subgroups Argyle et al. (2023); Chang et al. (2025).
However, growing evidence challenges these optimistic claims. LLM-generated synthetic opinions reveal fundamental limitations of prompting-based alignment. While persona prompting approximates US opinion distributions, performance degrades substantially for non-English populations Argyle et al. (2023); Ma et al. (2024); Münker (2025b, a). This suggests prompting privileges cultural contexts embedded in training data rather than reliably grounding outputs in external reality.
Moreover, prompting produces unstable representations. Minor prompt variations dramatically affect behavioral outputs Ma et al. (2024), and RLHF training compounds this instability Tjuatja et al. (2024). These findings indicate that prompting inadequately controls model behavior, failing to consistently activate intended knowledge representations.
2.3 Methodological and Ethical Concerns
Beyond technical limitations, using LLMs as human proxies raises serious methodological questions. Training data biases can systematically misrepresent marginalized groups Abid et al. (2021); Hutchinson et al. (2020). The disembodied nature of LLMs means they lack grounding in physical reality, cultural contexts, physical environments, and interpersonal relationships—that fundamentally shapes human cognition and decision-making Hussein (2012); Argyle et al. (2023).
Ethically, deploying LLMs as participant replacements raises concerns about informed consent, potential misuse, and oversimplification of complex human behaviors. These concerns intensify when researchers make causal claims based on simulated data without validating empirical realism.
Our Position
We argue that LLM-based simulation can contribute meaningfully to social science, but only when coupled with rigorous validation against real-world data. Claims about simulation validity must be empirically demonstrated, not assumed. This paper provides a framework for such validation.
@@ 3 Modeling Online Social Networks
We present a formal framework for Twins of Online Social Networks (TWONs) that captures both user behavior (Sect. 3.1) and platform mechanics (Sect. 3.2). This framework enables systematic comparison between simulated and observed behaviors. Figure 1 illustrates key concepts.
Figure 1: Illustrating example: Agent ($a^{2}$) sends a message to agent ($a^{1}$), triggering a model update and reply generation. Agent ($a^{2}$) receives a curated list of incoming messages at $t+1$ and updates its model accordingly. This illustrates the interplay between agent behavior ($b$) and network mechanics ($r$).
@@ 3.1 Modeling Agents
$M$ is a set of messages, where $m_{t}^{(i,j)}$ is a message from agent $i$ to agent $j$ at time $t$. Consequently, $m_{t}^{(i,\cdot)}$ indicates a message that agent $i$ transmits to all other agents and $M_{t}^{(\cdot,i)}$ indicates all messages sent to agent $i$. Accordingly, $M_{t}^{i}$ indicates all messages that agent $i$ sends and receives at time $t$.
@@
The communicative behavior $b^{i}(\cdot)$ of the agent $i$ is a function that maps its current state $a_{t}^{i}$ to the messages $m_{t+1}^{(i,j)}$ he is sending next to agents $j$: $b^{i}_{t}(a_{t}^{i})=M_{t}^{(i,j)}$.
@@ 3.2 Modeling Network Mechanics
The mechanics of the communication channel (here, an online social network) is defined by a function $r^{i}_{t}(\cdot)$ that adapts the agent $i$’ to the incoming messages $M_{t}^{(\cdot,i)}$ to $\widetilde{M}_{t}^{(\cdot,i)}$, returning a manipulated list of messages. Standard manipulations of the list are filtering or ranking, but could also include the manipulation of content or the addition of messages to the list that have not been directly sent between agents. Note that this function can be personalized and be adapted over time.
@@
4 Simulating TWONs
Based on the formal framework introduced in Sect. 3 we define the required tasks to construct a TWON by replicating its behavior at the user (Sect. 4) and system level (Sect. 4). Finally, this allows us to check the empirical realism of such TWONs (Sect. 4).
Behavior $b$ Agent $a^{i}$ Task Type Input $\widetilde{M}_{t}^{(\cdot,i)}$ Output $m_{t+1}^{(i,\cdot)}$ Loss $L_{b}$ Metric(s) $q$
@@ 4 Simulating TWONs
The task of machine learning an agent-based simulation of human social media communication behavior at user level is to estimate the function $b^{i}(\cdot)$ of each agent $i$, which given real-world observations of discourses $(M_{0}^{i},M_{1}^{i},...,M_{t}^{i})$ predicts its next message $m_{t+1}^{i}$. Thus, the objective can be formulated as minimizing a loss function
@@
that compares the predicted $\widehat{M}_{t}^{(\cdot,i)}$ with its observed value.
@@ 5 Case Study Design: Imitating $\mathbb{X}$ Users
We now instantiate our framework using data from $\mathbb{X}$. Our implementation focuses on user behavior function $b^{i}$ while acknowledging simplifications in other framework components. We estimate communicative behavior by minimizing $L_{b}(\widehat{m}_{t+1},m_{t+1})$ across three action types: posts, replies, and reply-likelihood.
@@ 6 Fitting Agents to User Data from $\mathbb{X}$
Our experiments are based on two datasets, English and German, collected from $\mathbb{X}$. The datasets are collected around keywords concerning the political discourses in the US and Germany during the first half of 2023. The samples contain two types of content: a) Tweets (posts) from delegates of the national parliament concerning political decisions – DE: $155.000$, EN: $930.000$ and b) replies from regular $\mathbb{X}$ users towards these decisions or opinions – DE: $185.000$, EN: $17.800.000$. In addition, we labeled the topics of posts and replies using zero-shot prompt classification with Llama 3.1 70B Zheng et al. (2023). This yields two distinct agent types, the a) political actor for posts and the b) unspecified $\mathbb{X}$ user for replies, both across two languages.
@@
Retweets: Tweets starting with RT mark a shared post. The author does not add new content, but replicates the original text and shares it with their followers. We exclude these samples because we cannot verify if the shared content is in line with the retweeting author’s opinion.
@@
We train two model adapters on top of Llama-3.2-3B-Instruct (Dubey et al., 2024) using the supervised fine-tuning (SFT) paradigm through the transformer reinforcement learning (TRL) library using the default pipeline. The trained adapters are unmodified LoRA (Hu et al., 2021) matrices provided by the parameter-efficient fine-tuning (PEFT) package. We optimize for two objectives (posting and replying) with contents from the original users as the labeled data.
@@ 7.1 Evaluation of Imitated Posting and Replying Behavior
An initial observation is that BLEU scores Papineni et al. (2002) in all in-context settings are low. This is also true for the n-gram metrics. However, the fine-tuning-based approach shows significant improvements specifically in the English reply task, but also for uni-gram matches and the overall length ratio. Beyond such token-overlap-metrics, we tested semantic similarity metrics: we embedded all samples using $\mathbb{X}$-fine-tuned BERT (Zhang et al., 2023) and calculated the distance between generations and original samples. Again, the results generated from the fine-tuned model show a significantly shorter distance to its original counterpart compared to the in-context model. The correlation between the predicted emotions in the original and generated samples using a BERT-based multi-label emotion classifier (Barbieri et al., 2020) is less clear. The fine-tuned approach only marginally outperforms the prompt-only approach on the posting task.
@@ 7.4 Imitation Limitations
Our experimental setup has further limitations regarding both data selection and the system’s capacity to generate discourses containing well-structured standpoints and arguments. Although our models demonstrate such capabilities under certain conditions, their performance characteristics primarily reflect the behavioral patterns of the most active users within our dataset, thus mirroring the distinctive communication dynamics observed in the selected $\mathbb{X}$ community. This sampling bias raises questions about the generalizability of our findings in different contexts of social networks and user populations.
@@ 8.2 Future Research
Unfortunately, our research objectives, demonstrating "empirical realism" (aka "ecological validity") of our simulation, do not allow us to use synthetic personas rather than real user data. Creating fictional user profiles with diverse characteristics would eliminate privacy concerns, but without demonstrating the validity of the simulation. We argue that you can only claim the applicability of findings produced via social simulations to the real world once you have proven the "empirical realism" (aka "ecological validity") of your simulation. The key component for showing this is replicability on an individual level, which requires real-world data from individuals. Synthetic personas cannot be used to prove "empirical realism", but real-world data is essential to prove the empirical realism of synthetic personas. And once you prove the empirical realism of synthetic personas with real-world data, you again have a sufficiently similar replica of real-world users. Thus, you cannot have both: A valid and provable realistic social simulation without using real-world user data to model agents and measure their empirical realism.